# An Equivalence Between Compatibility and Deterministic Underlying States in Quantum Mechanics

### v1

@@ Abstract
I Introduction
One of the most debated fundamental problems in modern physics is whether the quantum formalism’s description of nature is complete, or if it could be supplemented with additional, or “hidden”, variables [1, 2, 3, 4, 5, 6, 7]. While hidden-variable models have been constructed which (partially or fully) reproduce the predictions of quantum mechanics [5, 8, 9, 7], various impossibility theorems, notably Bell’s [10, 11] and Kochen and Specker’s [4], have shown that such attempts inevitably violate certain consistency conditions which we would arguably want from a successful description of reality [12, 13, 10, 4]. These theorems differ in the conditions they impose, but a common feature is their indirect dependence on incompatible observables [4, 11, 14, 15].
@@ I Introduction
To begin with, we argue that, contrary to the typical description of updated states as post-measurement states, updated states are theoretically constructed as pre-measurement states, with the qualifier post-measurement state regarded instead as a merely circumstantial attribute. That is, instead of being updated by propositions in the simple past (“a measurement of $\hat{A}$ was made, and the value $\alpha$ was obtained”), we argue that quantum states are theoretically updated to ensure the validity of propositions in the first conditional (“if a measurement is made of $\hat{A}$, the result will be found to be $\alpha$”) [24]. This approach aligns with multiple perspectives on quantum mechanics [25, 26, 27, 28, 29]. By this account, we argue Lüders’ rule should translate into underlying-state models as conditional probability, because considerations about interactions with measuring apparatuses are automatically removed. Using Bayes’ rule and Kolmogorov’s extension theorem [30], we then show that a set of quantum observables admits a state-updating underlying-state model (where states update via conditional probability) if and only if that set consists of pairwise compatible observables. A single pair of incompatible observables is sufficient to rule out a scenario being representable by a state-updating hidden-variable model. Building on this result, we analyse ontological and metaphysical views typically associated with the absence of hidden variables in quantum mechanics.
II Quantum state update
We first argue that updated states in quantum mechanics are constructed as pre-measurement states, with the qualifier post-measurement state, typically assumed to be their defining feature, instead being merely a circumstantial attribute. We show that this corresponds to an objective description of Lüders’ rule. Note that this is an account of the theoretical quantum update mechanism taken as an axiom in standard quantum mechanics. It should not be misunderstood as a particular interpretation of any physical process which occurs when quantum systems are measured in actual experiments.
@@ II Quantum state update
III Order-dependent predictions and incompatibility
We will denote the distribution of an observable $\hat{B}$ in the state $T_{[\hat{A}=\alpha]}(\psi)$ by $P_{\psi}[\hat{B}=\cdot\ |\hat{A}=\alpha]$. That is, for each $\beta\in\sigma(\hat{B})$ we have
@@ III Order-dependent predictions and incompatibility
IV State-updating Deterministic underlying-state models
Deterministic underlying-state models, as often used in the contextuality and nonlocality literature [17, 32, 18], are defined as follows:
@@ IV State-updating Deterministic underlying-state models
V Incompatibility obstructs state-updating underlying-state models
We now show that incompatibility obstructs the existence of state-updating underlying-state models. If a deterministic state-updating underlying-state model exists for a set of observables $\mathcal{O}_{S}$, then for any pair of observables $A,B\in\mathcal{O}_{S}$, any pure state $\psi$ and any values $\alpha\in\sigma(A)$, $\beta\in\sigma(B)$,
@@ V Incompatibility obstructs state-updating underlying-state models
VI Discussion
If our considerations regarding the quantum state update are correct, Proposition 1 provides a straightforward method for ruling out (deterministic) underlying-state models of quantum mechanics. The proof is independent of specific states and observables, as well as space-time considerations; it establishes a direct connection between compatibility and underlying states, demonstrating that incompatibility is sufficient to obstruct such models.
@@ VI Discussion
(clearly, $P_{\psi}[\hat{A}_{i}=\cdot\ |\hat{A}_{1}=\alpha_{1},\dots,\hat{A}_{i-1}=\alpha_{i-1}]$ denotes the distribution of $\hat{A}_{i}$ in the state $(T_{[\hat{A}_{i-1}=\alpha_{i-1}]}\circ\dots\circ T_{[\hat{A}_{1}=\alpha_{1}]})(\psi)$). Eq. (17) follow from straightforward calculations, as shown e.g., in Ref. [20]. In particular, it follows from our analysis of the state update mechanism that contextual deterministic underlying-state models are inconsistent with the canonical postulates of quantum mechanics (which include the collapse postulate).
@@
Acknowledgements.

### v2

@@ Abstract
Introduction— One of the most debated fundamental problems in modern physics is whether the quantum formalism’s description of nature is complete, or if it could be supplemented with additional, or “hidden”, variables [1, 2, 3, 4, 5, 6, 7]. While hidden-variable models have been constructed which (partially or fully) reproduce the predictions of quantum mechanics [5, 8, 9, 7], various impossibility theorems, notably Bell’s [10, 11] and Kochen and Specker’s [4], have shown that such attempts inevitably violate certain consistency conditions which we would arguably want from a successful description of reality [12, 13, 10, 4, 14]. These theorems differ in the conditions they impose, but a common feature is their indirect dependence on incompatible observables [4, 11, 15, 16].
@@
Fine [25, 26] and Malley [14, 27] have previously argued that incompatibility is sufficient to rule out underlying states. Our work differ from theirs in the consistency conditions we pose: conditional probability (our main assumption) is a theorem in Malley’s work [14], while their consistency conditions follows from ours.
This paper is structured as follows. To begin with, we argue that, contrary to the typical description of updated states as post-measurement states, updated states are theoretically constructed as pre-measurement states, with the qualifier post-measurement state regarded instead as a merely circumstantial attribute. That is, instead of being updated by propositions in the simple past (“a measurement of $\hat{A}$ was made, and the value $\alpha$ was obtained”), we argue that quantum states are theoretically updated to ensure the validity of propositions in the first conditional (“if a measurement is made of $\hat{A}$, the result will be found to be $\alpha$”) [28]. This approach aligns with multiple perspectives on quantum mechanics [29, 30, 31, 32, 33]. By this account, we argue Lüders’ rule should translate into underlying-state models as conditional probability, because considerations about interactions with measuring apparatuses are automatically removed. Using Bayes’ rule and Kolmogorov’s extension theorem [34], we then show that a set of quantum observables admits a state-updating underlying-state model (where states update via conditional probability) if and only if that set consists of pairwise compatible observables. A single pair of incompatible observables is sufficient to rule out a scenario being representable by a state-updating hidden-variable model. Building on this result, we analyse ontological and metaphysical views typically associated with the absence of hidden variables in quantum mechanics.
Quantum state update— We first argue that updated states in quantum mechanics are constructed as pre-measurement states, with the qualifier post-measurement state, typically assumed to be their defining feature, instead being merely a circumstantial attribute. We show that this corresponds to an objective description of Lüders’ rule. Note that this is an account of the theoretical quantum update mechanism taken as an axiom in standard quantum mechanics. It should not be misunderstood as a particular interpretation of any physical process which occurs when quantum systems are measured in actual experiments.
@@
Order-dependent predictions and incompatibility— We will denote the distribution of an observable $\hat{B}$ in the state $T_{[\hat{A}=\alpha]}(\psi)$ by $P_{\psi}[\hat{B}=\cdot\ |\hat{A}=\alpha]$. That is, for each $\beta\in\sigma(\hat{B})$ we have
@@
State-updating Deterministic underlying-state models— Deterministic underlying-state models, as often used in the contextuality and nonlocality literature [18, 36, 19], are defined as follows:
@@
Incompatibility obstructs state-updating underlying-state models— We now show that incompatibility obstructs the existence of state-updating underlying-state models. If a deterministic state-updating underlying-state model exists for a set of observables $\mathcal{O}_{S}$, then for any pair of observables $A,B\in\mathcal{O}_{S}$, any pure state $\psi$ and any values $\alpha\in\sigma(A)$, $\beta\in\sigma(B)$,
@@
Discussion— If our considerations regarding the quantum state update are correct, Proposition 1 provides a straightforward method for ruling out (deterministic) underlying-state models of quantum mechanics. The proof, similar to those in Refs. [14, 27], is independent of specific states and observables, as well as space-time considerations; it establishes a direct connection between compatibility and underlying states, demonstrating that incompatibility is sufficient to obstruct such models.
@@
(clearly, $P_{\psi}[\hat{A}_{i}=\cdot\ |\hat{A}_{1}=\alpha_{1},\dots,\hat{A}_{i-1}=\alpha_{i-1}]$ denotes the distribution of $\hat{A}_{i}$ in the state $(T_{[\hat{A}_{i-1}=\alpha_{i-1}]}\circ\dots\circ T_{[\hat{A}_{1}=\alpha_{1}]})(\psi)$). Eq. (17) follow from straightforward calculations, as shown e.g., in Ref. [22] (in which it is also proved that state-updating models satisfy Fine’s consistency condition [26]). In particular, it follows from our analysis of the state update mechanism that contextual deterministic underlying-state models are inconsistent with the canonical postulates of quantum mechanics (which include the collapse postulate).
@@
Acknowledgments—At thanks Rafael Wagner for discussions.